When 1 mol of caffeine (C

8

H

10

N

4

O

2

) is burned in air, 4960 kJ of heat is evolved. Five

grams of caffei

ne is burned in a bomb calorimeter. The temperature is observed to

increase by 11.37°C. What is the heat capacity of the calorimeter?



1
Expert's answer
2021-10-28T03:00:26-0400

1 mol of caffeine weighs 194.2 g (its molar mass). Find the amount of heat evolved when 5 g of caffeine is burned:

"Q = 4960 * 5\/194.2 = 127.7 (kJ)"

Find the heat capacity of the calorimeter (the quantity of heat absorbed by the calorimeter for each 1°C rise in temperature):

"C = Q\/T = 127.7\/11.37 = 11.23 (kJ\/" °C)

Answer: 11.23 kJ/°C -- the heat capacity of the calorimeter.